Hello I would like to have help on finding the solution of two simultaneous equations.
4x-3=15y
-4×+12=-6y

Also, I would like to know how I can make an equation of of two given points (4,-3) and (5,-1)
Thank you.

Answer by stanbon(75887)

finding the solution of two simultaneous equations.
4x-3=15y
-4×+12=-6y
------
Add to get:
9 = 9y
y = 1
========
Solve for "x":
4x-3 = 15y
4x-3 = 15
4x = 18
x = 9/2
======================
Also, I would like to know how I can make an equation of of two given points (4,-3) and (5,-1)
slope = (-1--3)/(5-4) = 2/1 = 2
-------
Form: y = mx + b
Using (4,-3) and m = 2, solve for "b":
-3 = 2*4 + b
b = -11
------
Equation:
y = 2x - 11
-------------------
Cheers,
Stan H.
===================